OpenStudy (anonymous):

For a certain river, suppose the drought length Y is the number of consecutive time intervals in which the water supply remains below a critical value y0 (a deficit), preceded by and followed by periods in which the supply exceeds this critical value (a surplus). An article proposes a geometric distribution with p = 0.369 for this random variable. (Round your answers to three decimal places.) (a) What is the probability that a drought lasts exactly 3 intervals? At most 3 intervals? (b) What is the probability that the length of a drought exceeds its mean value by at least one standard deviation?

OpenStudy (anonymous):

That's wrong =/ I have the solutions but they don't have the steps. I'm just trying to understand how they got there =/

OpenStudy (anonymous):

a) exactly 3 P( X = 3) = (1- .369)^2 * .369 = .14692 at most 3 P ( X <= 3 ) = P(X = 1) + P( X = 2) + P( X = 3) = .369 + (1-.369)(.369) + (1- .369)^2 * .369 =

OpenStudy (anonymous):

They should be 0.093 and 0.841 :/

OpenStudy (anonymous):

you sure you copied the problem correctly and the solutions

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

you want the drought to last 3 intervals. so you need to top off the 3 drought intervals with a surplus interval. label a drought interval with F. label surplus interval by S P( FFFS) = ( 1- .369) ^3 (.369)

OpenStudy (anonymous):

Ahh okay that makes more sense

OpenStudy (anonymous):

the distribution we are using is P( X = k) = (1-p)^k*p , k = 0,1,2,3, ... a) exactly 3 P( X = 3) = (1- .369)^3 * .369 b) at most 3 P ( X <= 3 )= P(X = 0) + P( X = 1) + P( X = 2) + P(X=3) =(1-.369)^0*(.369) + (1-.369)^1*(.369) + (1-.369)^2*(.369) + (1-.369)^3*(.369) = .841467

OpenStudy (anonymous):

Okay, I understand that so far

OpenStudy (anonymous):

you dont need to memorize the distribution formula we know the success occurs on the fourth trial, since we want 3 consecutive drought intervals. ie P( FFFS)

OpenStudy (anonymous):

Okay so what about b?

OpenStudy (anonymous):

mean = ( 1- p) / p = (1-.369) / .369 = 1.71 st. dev = sqrt[ ( 1- p) / p^2] = 2.1527 P ( Y >= 1.711 +2.1527) = P( Y >= 3.8627 ) = P( Y > 4)

OpenStudy (anonymous):

P(Y > 4) = 1 - P( Y <= 4) or we can save time using the formula: P( Y >= k ) = (1 - p)^k P( Y > 4) = P( Y >= 5) = ( 1- .369)^5 = .1000

OpenStudy (anonymous):

The answer they gave is 0.159